OpenStudy (anonymous):

consider the region bounded by y=1/x,x=1 and x=M(m is a real number larger then 1) Calculate the volume of the solid obtained by rotating the region about the x-axis. whaT HAPPENS to the volume of the solid when M gets larger?

OpenStudy (anonymous):

HAY

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

consider the region bounded by y=1/x,x=1 and x=M(m is a real number larger then 1) Calculate the volume of the solid obtained by rotating the region about the x-axis. whaT HAPPENS to the volume of the solid when M gets larger?

myininaya (myininaya):

So this is the picture we have before the rotation: 

myininaya (myininaya):